### v1

# A Characterization of Metaplectic Time-Frequency Representations

Abstract.
We characterize all time-frequency representations that satisfy a general covariance property: any weak*-continuous bilinear mapping that intertwines time-frequency shifts on the configuration space with time-frequency shifts on phase space is a multiple of a metaplectic time-frequency representation.
Key words and phrases:
Time-frequency representations, metaplectic operators, symplectic group, Schwartz kernel theorem.
2020 Mathematics Subject Classification:
81S30, 22E46
I. Shafkulovska was funded in part by the Austrian Science Fund (FWF) [10.55776/P33217] and [10.55776/Y1199]. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
1. Introduction
Many important time-frequency representations possess the covariance property: a time-frequency shift of a function is transformed to a shift of the time-frequency representation. For one of the most central time-frequency representations, namely the Wigner distribution, this property looks as follows. Let
[…]
be the (cross-) Wigner distribution of two functions $f,g\in L^{2}({\mathbb{R}}^{d})$. The translation operator or shift on $L^{2}({\mathbb{R}}^{d})$ is defined by $T_{x}f(t)=f(t-x)$ for $t,x,\omega\in{\mathbb{R}}^{d}$, and the frequency shift or modulation operator is given by $M_{\omega}f(t)=e^{2\pi i\omega\cdot t}f(t)$. Its Fourier transform is $\widehat{M_{\omega}f}=T_{-\omega}\hat{f}$, whence frequency shift. Writing $\lambda=(x,\omega)$ for a point in the time-frequency space, the symmetrized composition is the time-frequency shift
$\rho(\lambda)=\rho(x,\omega)=T_{x/2}M_{\omega}T_{x/2},\quad\lambda=(x,\omega)\in{{\mathbb{R}}^{2d}}\,.$
Then the covariance property states that for all $\lambda,z\in{{\mathbb{R}}^{2d}},f,g\in L^{2}({\mathbb{R}}^{d})$
(1.2) $W(\rho(\lambda)f,\rho(\lambda)g)(z)=W(f,g)(z-\lambda)=(T_{\lambda}W(f,g))(z)\,.$
Many other useful time-frequency representations possess this useful and intuitive property, for instance, the spectrogram or the Born-Jordan distribution. Therefore, covariant time-frequency representations have been studied intensely in the engineering literature and in physics, and ultimately were completely classified by L. Cohen [2, 3]. If a time-frequency representation ${\mathfrak{R}}(f,g)$ is covariant, then ${\mathfrak{R}}$ must have the form ${\mathfrak{R}}(f,g)=W(f,g)\ast\Psi$ for some distribution $\Psi\in{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$. See also [13, Sec. 4.5] for a mathematical treatment of Cohen’s class.
A second look at the Wigner distribution reveals a more delicate structure. By viewing the Wigner distribution as a sequilinear mapping, we can apply different time-frequency shifts to $f$ and $g$ and obtain the following formula (e.g., [10, Prop. 1.94] or [13, Prop. 4.3.2])
$W(\rho(\lambda)f,\rho(\mu)g)(z)$ $=e^{\pi i\lambda\cdot{\mathcal{J}}\mu}e^{2\pi iz\cdot{\mathcal{J}}(\lambda-\mu)}W(f,g)\big{(}z-\frac{\lambda+\mu}{2}\big{)}$
(1.3) $=\rho\Big{(}\tfrac{\lambda+\mu}{2},{\mathcal{J}}(\lambda-\mu)\Big{)}W(f,g)(z)\,,$
with ${\mathcal{J}}=\begin{pmatrix}0&I_{d}\\ -I_{d}&0\end{pmatrix}$. Thus the Wigner distribution with two different time-frequency shifts is not only a shift of the Wigner distribution, but also contains a modulation, in other words, the result is a time-frequency shift of the original Wigner distribution. Formula (1.3) is a general covariance property, and it says that the transform of time-frequency shifts is again a time-frequency shift. Note, however, that on the right-hand side the time-frequency shift has twice as many variables. The general covariance property (1.3) is a fundamental property of the Wigner distribution and plays an important role in the time-frequency approach to pseudodifferential operators, see e.g.,  [14] and the recent monograph [4].
We say that a time-frequency representation, i.e., a bilinear mapping ${\mathfrak{R}}:{\mathcal{S}({{\mathbb{R}}^{d}})}\times{\mathcal{S}({{\mathbb{R}}^{d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$, possesses the general covariance property, if it maps time-frequency shifts of $f,g\in{\mathcal{S}({{\mathbb{R}}^{d}})}$ to a time-frequency shift of ${\mathfrak{R}}(f,g)$. Formally, there exists a function $\Phi_{{\mathfrak{R}}}:{\mathbb{R}}^{4d}\to{\mathbb{R}}^{4d}$, such that for all $f,g\in{\mathcal{S}({{\mathbb{R}}^{d}})}$ and all $\lambda,\mu\in{{\mathbb{R}}^{2d}}$ one has
(1.4) ${\mathfrak{R}}(\rho(\lambda)f,\rho(\gamma)g)=c\rho(\Phi_{\mathfrak{R}}(\lambda,\gamma)){\mathfrak{R}}(f,g).$
Note that in this definition $\rho$ occurs with different dimensions. On the left-hand side $\rho(\lambda)f$ acts on a function of $d$ variables, on the right hand side $\rho(\Phi_{{\mathfrak{R}}}(\lambda,\mu))$ acts on a distribution on ${\mathbb{R}}^{2d}$. In the sequel we will use only the latter version, therefore, we will not introduce a different notation.
It is natural to assume that $\Phi$ is measurable. In this paper, we derive a complete classification of all time-frequency representations that satisfy the general covariance property (1.4).
For the formulation of the main result, we recall the definition of a metaplectic operator. By the theorem of Stone-von Neumann [10] for every symplectic matrix $\mathcal{A}\in{\mathrm{Sp}(2d,{\mathbb{R}})}$ there exists a unitary operator $\hat{\mathcal{A}}$ on $L^{2}({\mathbb{R}}^{d})$, such that
(1.5) $\rho(\mathcal{A}\lambda)=\hat{\mathcal{A}}\rho(\lambda)\hat{\mathcal{A}}^{-1},$
that is, $\hat{\mathcal{A}}$ arises as an intertwining operator, and is called the metaplectic operator associated to $\mathcal{A}$.
With these preparations, the main result about time-frequency representations obeying the general covariance property goes as follows.
Theorem 1.1 (Main result).
Let ${\mathfrak{R}}:{\mathcal{S}({{\mathbb{R}}^{d}})}\times{\mathcal{S}({{\mathbb{R}}^{d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ be a non-zero, bilinear, separately weak*-continuous mapping satisfying the general covariance property  (1.4) with a measurable function $\Phi$. Then there exist $a\in{\mathbb{C}}$, $a\neq 0$, and $\mathcal{A}\in{\mathrm{Sp}(4d,{\mathbb{R}})}$ such that
(1.6) ${\mathfrak{R}}(f,g)=a\,\hat{\mathcal{A}}(f\otimes g),\qquad f,g\in{\mathcal{S}({{\mathbb{R}}^{d}})}.$
The general covariance property seems to be very strong. Although we start with a mapping from ${\mathcal{S}({{\mathbb{R}}^{d}})}\times{\mathcal{S}({{\mathbb{R}}^{d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$, the actual mapping properties are much stronger.
Corollary 1.2.
Let ${\mathfrak{R}}:{\mathcal{S}({{\mathbb{R}}^{d}})}\times{\mathcal{S}({{\mathbb{R}}^{d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$. Under the assumptions of Theorem 1.1, ${\mathfrak{R}}$ has the following properties.
  1. (a)
    The function $\Phi_{\mathfrak{R}}$ is unique and linear.
  2. (b)
    ${\mathfrak{R}}$ maps ${\mathcal{S}({{\mathbb{R}}^{d}})}\times{\mathcal{S}({{\mathbb{R}}^{d}})}$ into ${\mathcal{S}({{\mathbb{R}}^{2d}})}$.
  3. (c)
    ${\mathfrak{R}}$ is non-degenerate.
  4. (d)
    ${\mathfrak{R}}$ extends to an isometry ${\mathfrak{R}}:{L^{2}({\mathbb{R}}^{d})}\times{L^{2}({\mathbb{R}}^{d})}\to{L^{2}({{\mathbb{R}}^{2d}})}$, i.e.,
    (1.7) $\lVert{\mathfrak{R}}(f,g)\rVert_{2}=|a|\,\lVert f\rVert_{2}\lVert g\rVert_{2},\quad f,g\in{L^{2}({\mathbb{R}}^{d})}.$
  5. (e)
    The range $\{{\mathfrak{R}}(f,g):f,g\in{\mathcal{S}({{\mathbb{R}}^{d}})}\}$ spans a dense subspace of ${\mathcal{S}({{\mathbb{R}}^{2d}})}$, of $L^{2}({\mathbb{R}}^{d})$, and a weak-dense subspace of ${\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$.
Note that we only assume that $\Phi_{\mathfrak{R}}$ is measurable!
It turns out that time-frequency representations of the form $\hat{\mathcal{A}}(f\otimes\bar{g})$ were recently introduced by Cordero and Rodino [8] as a generalization of the Wigner distribution, the short-time Fourier transform, or Beyer’s time-frequency representations [1]. They are called metaplectic time-frequency representations or the metaplectic Wigner distributions associated to $\mathcal{A}$, or shortly, the $\mathcal{A}$-Wigner distributions. In a series of papers, Cordero, Rodino, and their collaborators  [7, 5, 6] have shown that metaplectic time-frequency representations are a versatile tool for the refined investigation of pseudodifferential operators and function spaces. Uncertainty principles for metaplectic time-frequency representations were investigated in [15, 16].
Theorem 1.1 provides an intrinsic, structural argument as to why metaplectic time-frequency representations are natural and are bound to appear in time-frequency analysis.
Remark 1.3.
Usually, a time-frequency representation is treated as a sesquilinear form rather than as a bilinear form. Theorem 1.1 can be equivalently restated in terms of sesquilinear forms by a tiny modification. If ${\mathfrak{R}}$ is bilinear, then the corresponding sesquilinear form ¯$\overline{{\mathfrak{R}}}(f,g)\coloneqq{\mathfrak{R}}(f,\overline{g})$ is non-zero, continuous, and satisfies
¯$\overline{{\mathfrak{R}}}(\rho(\lambda)f,\rho(\gamma)g)=c\rho(\Phi_{\overline{\mathfrak{R}}}(\lambda,\gamma))\overline{{\mathfrak{R}}}(f,g),\quad\Phi_{\overline{{\mathfrak{R}}}}(\lambda_{1},\lambda_{2},\gamma_{1},\gamma_{2})=\Phi_{{{\mathfrak{R}}}}(\lambda_{1},\lambda_{2},\gamma_{1},-\gamma_{2}).$
The change of the intertwining function is due to $\overline{\rho(\gamma_{1},\gamma_{2})g}=\rho(\gamma_{1},-\gamma_{2})\overline{g}$. The conclusion of Theorem 1.1 ¯$\overline{{\mathfrak{R}}}(f,g)=a\,\hat{\mathcal{A}}(f\otimes\bar{g})$ remains the same.
The remainder of the paper is devoted to the proof of Theorem 1.1. We first use the continuity assumptions of ${\mathfrak{R}}$ to identify it with a continuous linear mapping ${\mathfrak{T}}:{\mathcal{S}({{\mathbb{R}}^{2d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ that satisfies a general covariance property. Here, the arguments are analytical and are based on a bilinear version of the Schwartz kernel theorem. In the next step, we exploit the underlying Lie group structure of the problem to deduce that $\Phi$ must be a linear mapping. Finally, the general covariance property (1.4) implies that $\Phi$ is symplectic. The arguments here are purely algebraic. We put these three aspects together to derive the final form of ${\mathfrak{R}}$.
We would like to offer our thanks and appreciation to Christian Bargetz whose advice helped us to compress our original proof of Lemma 3.1 to a few lines.
2. Symplectic matrices and metaplectic operators
The symmetric time-frequency shifts
$\rho(\lambda)=\rho(x,\omega)=T_{x/2}M_{\omega}T_{x/2},\quad\lambda=(x,\omega)\in{{\mathbb{R}}^{2d}},$
are a projective unitary representation of ${{\mathbb{R}}^{2d}}$. They satisfy
(2.1) $\rho(\lambda_{1}+\lambda_{2})=e^{\pi i\lambda_{1}\cdot{\mathcal{J}}\lambda_{2}}\rho(\lambda_{1})\rho(\lambda_{2}),\qquad{\mathcal{J}}=\begin{pmatrix}0&I_{d}\\ -I_{d}&0\end{pmatrix}.$
In particular, for all $n\in{\mathbb{N}}$ and $\lambda\in{{\mathbb{R}}^{d}}$ holds
(2.2) $\rho(\lambda)^{-1}=\rho(-\lambda),\qquad\text{ and }\quad\rho(n\lambda)=\rho(\lambda)^{n}$
The symplectic matrices are those matrices which preserve the standard symplectic form $[\lambda_{1},\lambda_{2}]=\lambda_{1}\cdot{\mathcal{J}}\lambda_{2}$, i.e.,
(2.3) ${\mathrm{Sp}(2d,{\mathbb{R}})}=\left\{\mathcal{A}\in{\mathbb{R}}^{2d\times 2d}:\mathcal{A}^{t}{\mathcal{J}}\mathcal{A}={\mathcal{J}}\right\}.$
This is a distinguished group of matrices in time-frequency analysis [13, 10, 9]. By the Stone-von Neumann theorem, for every symplectic matrix $\mathcal{A}\in{\mathbb{R}}^{2d\times 2d}$, there exists a unitary operator $\mathcal{A}$ that is unique up to scaling with $c\in{\mathbb{T}}$ [10, 13, 9] satisfying
(2.4) $\rho(\mathcal{A}\lambda)=\hat{\mathcal{A}}\rho(\lambda)\hat{\mathcal{A}}^{-1}\,.$
Thus $\hat{\mathcal{A}}$ arises as an intertwining operator, and is called the metaplectic operator associated to $\mathcal{A}$. The metaplectic operators are isomorphisms on the Schwartz space ${\mathcal{S}({{\mathbb{R}}^{d}})}$, the space of rapidly decaying functions [10, 9, 13], and by duality, on its dual space ${\mathcal{S}^{\prime}({{\mathbb{R}}^{d}})}$.
3. Analytic consequences
In this section, we exploit the weak*-continuity of the bilinear form ${\mathfrak{R}}$. Explicitly this means that for every sequence $(f_{n})_{n\in{\mathbb{N}}}\in{\mathcal{S}({{\mathbb{R}}^{d}})}$ converging to $f\in{\mathcal{S}({{\mathbb{R}}^{d}})}$ in the topology of ${\mathcal{S}({{\mathbb{R}}^{d}})}$ and every $g\in{\mathcal{S}({{\mathbb{R}}^{d}})}$, we have
(3.1) $\lim\limits_{n\to\infty}{\mathfrak{R}}(f_{n},g)={\mathfrak{R}}(f,g)\quad\text{in the weak*-topology on }{\mathcal{S}^{\prime}({{\mathbb{R}}^{d}})},$
and, likewise, for all $(g_{n})_{n\in{\mathbb{N}}}\in{\mathcal{S}({{\mathbb{R}}^{d}})}$ converging to $g\in{\mathcal{S}({{\mathbb{R}}^{d}})}$ and all $f\in{\mathcal{S}({{\mathbb{R}}^{d}})}$
(3.2) $\lim\limits_{n\to\infty}{\mathfrak{R}}(f,g_{n})={\mathfrak{R}}(f,g)\quad\text{in the weak*-topology on }{\mathcal{S}^{\prime}({{\mathbb{R}}^{d}})}.$
Throughout, we denote with $\langle T,f\rangle=T(f)$ the evaluation of the tempered distribution $T\in{\mathcal{S}^{\prime}({{\mathbb{R}}^{d}})}$ at $f\in{\mathcal{S}({{\mathbb{R}}^{d}})}$. We will use the following bilinear version of the Schwartz kernel theorem [20, p. 531].
Lemma 3.1.
For every separately weak*-continuous bilinear mapping
${\mathfrak{R}}:{\mathcal{S}({{\mathbb{R}}^{d}})}\times{\mathcal{S}({{\mathbb{R}}^{d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})},$
there exists a unique weak*-continuous linear operator ${\mathfrak{T}}:{\mathcal{S}({{\mathbb{R}}^{2d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ such that for all $f,g\in{\mathcal{S}({{\mathbb{R}}^{d}})}$ and all $H\in{\mathcal{S}({{\mathbb{R}}^{2d}})}$
(3.3) $\langle{\mathfrak{R}}(f,g),H\rangle=\langle{\mathfrak{T}}(f\otimes g),H\rangle.$
In addition, there exists a kernel $K=K_{\mathfrak{R}}\in\mathcal{S}^{\prime}({{\mathbb{R}}^{4d}})$ such that
(3.4) $\langle{\mathfrak{R}}(f,g),H\rangle=\langle K,f\otimes g\otimes H\rangle,\quad f,g\in{\mathcal{S}({{\mathbb{R}}^{d}})},\,H\in{\mathcal{S}({{\mathbb{R}}^{2d}})}.$
Proof.
As we are unaware of an explicit reference, we offer a proof sketch for completeness. The proof uses the same arguments as the numerous kernel theorems in functional analysis. We recall that ${\mathcal{S}({{\mathbb{R}}^{d}})}$ is a nuclear Fréchet space and hence a Montel space, i.e., bounded, closed sets in ${\mathcal{S}({{\mathbb{R}}^{d}})}$ are compact and vice versa, see [20, Prop. 34.4]. This property implies that a weak-continuous map from ${\mathcal{S}({{\mathbb{R}}^{d}})}$ to ${\mathcal{S}({{\mathbb{R}}^{d}})}^{\prime}$ is automatically continuous, when ${\mathcal{S}({{\mathbb{R}}^{d}})}^{\prime}$ is endowed with the strong topology (topology of uniform convergence on bounded sets) [20, p. 358]. Consequently, for fixed $g\in{\mathcal{S}({{\mathbb{R}}^{d}})}$ the map $f\in{\mathcal{S}({{\mathbb{R}}^{d}})}\mapsto{\mathfrak{T}}(f,g)\in{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ is strongly continuous, and likewise $g\in{\mathcal{S}({{\mathbb{R}}^{d}})}\mapsto{\mathfrak{T}}(f,g)\in{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ is strongly continuous for every $f\in{\mathcal{S}({{\mathbb{R}}^{d}})}$. Since ${\mathcal{S}({{\mathbb{R}}^{d}})}$ is a Fréchet space, the separate continuity implies the joint continuity of ${\mathfrak{T}}$ on the Cartesian product ${\mathcal{S}({{\mathbb{R}}^{d}})}\times{\mathcal{S}({{\mathbb{R}}^{d}})}$ [18, Thm. 2.17].
Now we extend ${\mathfrak{T}}$ to a linear map on the algebraic tensor product ${\mathcal{S}({{\mathbb{R}}^{d}})}\times{\mathcal{S}({{\mathbb{R}}^{d}})}$. We endow ${\mathcal{S}({{\mathbb{R}}^{d}})}\times{\mathcal{S}({{\mathbb{R}}^{d}})}$ with the projective tensor product topology. This is by definition the finest topology such that the map $(f,g)\mapsto f\otimes g$ is continuous. As a consequence, ${\mathfrak{T}}$ is (strongly) continuous on ${\mathcal{S}({{\mathbb{R}}^{d}})}\times{\mathcal{S}({{\mathbb{R}}^{d}})}$. The completion of ${\mathcal{S}({{\mathbb{R}}^{d}})}\times{\mathcal{S}({{\mathbb{R}}^{d}})}$ with respect to this topology is the projective tensor product ${\mathcal{S}({{\mathbb{R}}^{d}})}\otimes_{\pi}\,{\mathcal{S}({{\mathbb{R}}^{d}})}$ and is canonically isomorphic with $\mathcal{S}({{\mathbb{R}}^{2d}})$ [20, Thm. 51.6.]. By continuity, ${\mathfrak{T}}$ can now be extended to a (strongly) continuous map from ${\mathcal{S}({{\mathbb{R}}^{d}})}\otimes_{\pi}\,{\mathcal{S}({{\mathbb{R}}^{d}})}\cong{\mathcal{S}({{\mathbb{R}}^{2d}})}$ to ${\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$. Now the Schwarz kernel theorem provides a kernel proving the representation (3.4). ∎
We now relate the bilinear kernel theorem to the setup in Theorem 1.1. We show that the linear extension ${\mathfrak{T}}$ of ${\mathfrak{R}}$ also satisfies a covariance property.
Lemma 3.2.
Let ${\mathfrak{R}}:{\mathcal{S}({{\mathbb{R}}^{d}})}\times{\mathcal{S}({{\mathbb{R}}^{d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ be a non-zero, separately weak*-continuous bilinear form satisfying the covariance property (1.4). Then there exists a non-zero weak*-continuous operator ${\mathfrak{T}}:{\mathcal{S}({{\mathbb{R}}^{2d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ and a function $\Phi:{{\mathbb{R}}^{4d}}\to{{\mathbb{R}}^{4d}}$ satisfying
(3.5) ${\mathfrak{T}}(f\otimes g)={\mathfrak{R}}(f,g),\quad f,g\in{\mathcal{S}({{\mathbb{R}}^{d}})}$
and the covariance property
(3.6) ${\mathfrak{T}}\rho(\lambda)F=c\rho(\Phi(\lambda)){\mathfrak{T}}F,\quad\lambda\in{{\mathbb{R}}^{4d}},\,F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}.$
The function $\Phi$ is measurable if and only if $\Phi_{\mathfrak{R}}$ is measurable.
Proof.
Since ${\mathfrak{R}}$ is separately weak-continuous, Lemma 3.1 asserts the existence of the associated linear operator ${\mathfrak{T}}$. We claim that
(3.7) $\Phi(\lambda)=\Phi(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})=\Phi_{\mathfrak{R}}(\lambda_{1},\lambda_{3},\lambda_{2},\lambda_{4}),\quad\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\in{{\mathbb{R}}^{d}}.$
has the desired property. The modification of $\Phi$ is a consequence of the reordering of time-frequency shifts for $\rho$ on ${{\mathbb{R}}^{2d}}\times{{\mathbb{R}}^{2d}}$ to ${\mathbb{R}}^{4d}$ as follows:
(3.8) $\rho(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})(f\otimes g)=\rho(\lambda_{1},\lambda_{3})f\otimes\rho(\lambda_{2},\lambda_{4})g,\quad\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\in{{\mathbb{R}}^{d}}.$
This implies that (3.6) is satisfied by the elementary tensors, and by linearity, on their linear span. Since time-frequency shifts are weak*-continuous on ${\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ [13, Cor. 11.2.22], the identity ${\mathfrak{T}}\rho(\lambda)=\rho(\Phi(\lambda)){\mathfrak{T}}$ extends from ${\mathcal{S}({{\mathbb{R}}^{d}})}\otimes{\mathcal{S}({{\mathbb{R}}^{d}})}$ to ${\mathcal{S}({{\mathbb{R}}^{2d}})}$ by continuity and density. ∎
4. Consequences of the General Covariance Property
In this section, we focus on the implications of the covariance property (1.4), and by extension, of the covariance property (3.6), and show that $\Phi$ is necessarily a symplectic linear transformation. For the proof we will use some basic, but deep facts about homomorphisms between Lie groups.
We split the proof into several lemmata.
Lemma 4.1.
Let ${\mathfrak{T}}:{\mathcal{S}({{\mathbb{R}}^{2d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ be a non-zero continuous linear operator satisfying the general covariance property (3.6). Then ${\mathfrak{T}}$ is injective.
Proof.
If ${\mathfrak{T}}F_{0}=0$ for a non-zero $F_{0}\in{\mathcal{S}({{\mathbb{R}}^{2d}})}$, then for all linear combinations $F=\sum_{j=1}^{k}c_{j}\rho(\lambda_{j})F_{0}$ we obtain
(4.1) ${\mathfrak{T}}F=\sum_{j=1}^{k}c_{j}{\mathfrak{T}}\rho(\lambda_{j})F_{0}=\sum_{j=1}^{k}c_{j}\rho(\Phi(\lambda_{j})){\mathfrak{T}}F_{0}=0.$
Therefore, ${\mathfrak{T}}$ vanishes on the dense subspace of finite linear combinations. By continuity, ${\mathfrak{T}}=0$, contradicting the assumption. ∎
Lemma 4.1 implies the no-degeneracy stated in Corollary 1.2 (c).
Recall that we denote the skew-symmetric form with $[\lambda,\mu]=\lambda\cdot{\mathcal{J}}\mu$, $\lambda,\mu\in{{\mathbb{R}}^{2d}}$.
Lemma 4.2.
Let ${\mathfrak{T}}:{\mathcal{S}({{\mathbb{R}}^{2d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ be a non-zero continuous linear operator satisfying the general covariance property (3.6). Then
(4.2) $\rho(\Phi(\lambda+\mu)){\mathfrak{T}}F=e^{\pi i([\lambda,\mu]-[\Phi(\lambda),\Phi(\mu)])}\rho(\Phi(\lambda)+\Phi(\mu)){\mathfrak{T}}F,\quad F\in{\mathcal{S}({{\mathbb{R}}^{2d}})},\ \lambda,\mu\in{{\mathbb{R}}^{4d}},$
and
(4.3) $[\lambda,\mu]-[\Phi(\lambda),\Phi(\mu)]\in{\mathbb{Z}},\quad\lambda,\mu\in{{\mathbb{R}}^{4d}}.$
Proof.
Let $F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}$ be a non-zero function, $\lambda,\mu\in{{\mathbb{R}}^{4d}}$. Then
(4.4) $\begin{split}\rho(\Phi(\lambda+\mu)){\mathfrak{T}}F&=c^{-1}\,{\mathfrak{T}}\rho(\lambda+\mu)F=c^{-1}\,e^{\pi i[\lambda,\mu]}{\mathfrak{T}}\rho(\lambda)\rho(\mu)F\\ &=c\,e^{\pi i[\lambda,\mu]}\rho(\Phi(\lambda))\rho(\Phi(\mu)){\mathfrak{T}}F\\ &=c\,e^{\pi i([\lambda,\mu]-[\Phi(\lambda),\Phi(\mu)])}\rho(\Phi(\lambda)+\Phi(\mu)){\mathfrak{T}}F\\ &=c\,e^{-\pi i([\lambda,\mu]-[\Phi(\lambda),\Phi(\mu)])}\rho(\Phi(\lambda)+\Phi(\mu)){\mathfrak{T}}F.\end{split}$
The last equality here is obtained by switching the places of $\lambda$ and $\mu$ (recall that $[\lambda,\mu]=-[\mu,\lambda]$). By Lemma 4.1, the chain of equalities can only hold if the phase factors are equal, i.e.,
(4.5) $e^{2\pi i([\lambda,\mu]-[\Phi(\lambda),\Phi(\mu)])}=1.$
This is equivalent to (4.3). ∎
To show that $\Phi$ is linear, we introduce the subgroup
(4.6) $H=\{\nu\in{{\mathbb{R}}^{4d}}:\ \exists c_{\nu}\in{\mathbb{C}}\text{ with }\rho(\nu){\mathfrak{T}}F=c_{\nu}{\mathfrak{T}}F\text{ for all }F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}\},$
This is the projective kernel of the representation $\rho$ restricted to the subspace $\mathrm{ran}\,{\mathfrak{T}}=\{{\mathfrak{T}}F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}:F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}\}$, and by continuity of ${\mathfrak{T}}$ and $\rho(\lambda)$, $\lambda\in{{\mathbb{R}}^{4d}}$, to its closure $\overline{\mathrm{ran}\,{\mathfrak{T}}}$ in ${\mathcal{S}({{\mathbb{R}}^{2d}})}$. Let
(4.7) $q_{H}:{{\mathbb{R}}^{4d}}\to{{\mathbb{R}}^{4d}}/H$
denote the quotient map.
Lemma 4.3.
Let ${\mathfrak{T}}:{\mathcal{S}({{\mathbb{R}}^{2d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ be a non-zero continuous linear operator satisfying the general covariance property (3.6). Further, let $H$ be the projective kernel of $\rho$ defined in (4.6). Then $H$ is a closed subgroup of ${{\mathbb{R}}^{4d}}$.
Proof.
Since $\rho$ is a projective representation of ${{\mathbb{R}}^{4d}}$, it follows that $H$ is a group:
(4.8) $\rho(\nu+\eta){\mathfrak{T}}F=e^{\pi i[\nu,\eta]}c_{\nu}c_{\eta}{\mathfrak{T}}F,\quad F\in{\mathcal{S}({{\mathbb{R}}^{2d}})},\,\nu,\eta\in H.$
The continuity of the mapping $\lambda\mapsto\rho(\lambda){\mathfrak{T}}F$ implies that $H$ is closed and the mapping $H\to{\mathbb{C}},\ \nu\mapsto c_{\nu},$ is continuous. To see this, let $(\nu_{n})_{n\in{\mathbb{N}}}\subseteq H$ be a convergent sequence with limit $\mu\in{{\mathbb{R}}^{4d}}$. Then for all $F,G\in{\mathcal{S}({{\mathbb{R}}^{2d}})}$
(4.9) $\langle\rho(\nu){\mathfrak{T}}F,G\rangle=\langle\lim\limits_{n\to\infty}\rho(\nu_{n}){\mathfrak{T}}F,G\rangle=\lim\limits_{n\to\infty}c_{\nu_{n}}\langle{\mathfrak{T}}F,G\rangle.$
Since the limit in the middle converges, so does the limit on the right-hand side. Since ${\mathfrak{T}}$ is not zero, there exist $F,G$ with $\langle{\mathfrak{T}}F,G\rangle\neq 0$. Thus, $c_{\nu}=\lim\limits_{n\to\infty}c_{\nu_{n}}$ is well-defined and satisfies $\rho(\nu){\mathfrak{T}}F=c_{\nu}{\mathfrak{T}}F$, $F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}$. ∎
Lemma 4.4.
Let ${\mathfrak{T}}:{\mathcal{S}({{\mathbb{R}}^{2d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ be a non-zero continuous linear operator satisfying the general covariance property (3.6). Further, let $H$ be the projective kernel of $\rho$ defined in (4.6). Then the map
(4.10) $q_{H}\circ\Phi:{{\mathbb{R}}^{4d}}\to{{\mathbb{R}}^{4d}}/H,\quad\lambda\mapsto(q_{H}\circ\Phi)(\lambda)$
is injective.
Proof.
We want to show that for all $\lambda,\mu\in{{\mathbb{R}}^{4d}}$ and $\nu\in H$
(4.11) $\Phi(\lambda)=\Phi(\mu)+\nu\quad\implies\lambda=\mu.$
In a single chain of inequalities, we have for all $F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}$
(4.12) ${\mathfrak{T}}(\rho(\lambda)F)$ $=c\,\rho(\Phi(\lambda)){\mathfrak{T}}F=c\,\rho(\Phi(\mu)+\nu){\mathfrak{T}}F$
(4.13) =(2.1)${=}c\,e^{\pi i[\Phi(\mu),\nu]}\rho(\Phi(\mu))\rho(\nu){\mathfrak{T}}F$
(4.14) $=c\,c_{\nu}\,e^{\pi i[\Phi(\mu),\nu]}\rho(\Phi(\mu)){\mathfrak{T}}F$
(4.15) $={\mathfrak{T}}(c\,c_{\nu}\,e^{\pi i[\Phi(\mu),\nu]}\rho(\mu)F).$
By Lemma 4.1,
(4.16) $\rho(\lambda)F=c\,c_{\nu}\,e^{\pi i[\Phi(\mu),\nu]}\rho(\mu)F,$
or equivalently,
(4.17) $\rho(\lambda-\mu)F=c\,c_{\nu}\,e^{\pi i([\Phi(\mu),\nu]+[\lambda,\mu])}F.$
Since (4.17) holds for all $F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}$, the time-frequency shift $\rho(\lambda-\mu)$ must be a multiple of the identity, and therefore $\lambda=\mu$. ∎
We now prove that $\Phi$ is an invertible linear map.
Proposition 4.5.
Let ${\mathfrak{T}}:{\mathcal{S}({{\mathbb{R}}^{2d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ be a non-zero continuous linear operator satisfying the covariance property (3.6) with a measurable $\Phi$. Then $\Phi$ is an invertible linear map, i.e., there exists a matrix $\mathcal{A}\in\mathrm{GL}(4d,{\mathbb{R}})$ such that $\Phi(\lambda)=\mathcal{A}\lambda$, $\lambda\in{{\mathbb{R}}^{4d}}$.
Proof.
Let $H$ be the projective kernel of $\rho$ defined in (4.6), and $q_{H}$ the associated quotient map.
An injective Lie group homomorphism. By Lemma 4.2, for all $\lambda,\mu\in{{\mathbb{R}}^{4d}}$
(4.18) $\Phi(\lambda+\mu)-\Phi(\lambda)-\Phi(\mu)\in H.$
Thus,
(4.19) $(q_{H}\circ\Phi)(\lambda+\mu)=(q_{H}\circ\Phi)(\lambda)+(q_{H}\circ\Phi)(\mu),$
that is, $\pi=q_{H}\circ\Phi$ is a group homomorphism. By Lemma 4.4, it is injective.
Since $q_{H}$ is continuous and $\Phi$ is measurable, $\pi$ is measurable. By [17, Thm. 22.18], this implies that $\pi$ is continuous. The continuity of $\pi$ implies that it is even smooth, i.e., $\pi$ is a Lie group homomorphism from ${{\mathbb{R}}^{4d}}$ to $G={{\mathbb{R}}^{4d}}/H$. It induces a Lie algebra homomorphism between the corresponding Lie algebras ${{\mathbb{R}}^{4d}}$ and $\mathfrak{g}={\mathbb{R}}^{k}$.
Dimension count. Since $G$ is a quotient of ${{\mathbb{R}}^{4d}}$ and a closed subgroup, $k\leq 4d$ [21, Cor. 2.7.4.]. On the other hand, $d\pi$ is injective by [21, Cor. 2.7.4.]. Thus, $4d\leq k$, i.e., $k=4d$. This implies that the simply connected covering group $\widetilde{G}$ of $G$ is an Abelian Lie group with dimension $4d$ and thus $\widetilde{G}={{\mathbb{R}}^{4d}}$ and $H$ must be a discrete subgroup of ${{\mathbb{R}}^{4d}}$. Lastly, since
(4.20) $d\pi:{{\mathbb{R}}^{4d}}\to{{\mathbb{R}}^{4d}}$
is injective, it is invertible.
Comparing homomorphisms. By [21, Thm. 2.7.5.], there exists a unique homomorphism $\eta:{{\mathbb{R}}^{4d}}\to\tilde{G}$ such that $d\eta=d\pi$. Since $d\pi$ is invertible, $\eta$ is an automorphism. This means that there exists an invertible matrix $\mathcal{A}\in\mathrm{GL}(4d,{\mathbb{R}})$ such that $\eta(\lambda)=\mathcal{A}\lambda$.
If we go back to the quotient space $G$, the composition $\psi=q_{H}\circ\eta:{{\mathbb{R}}^{4d}}\to G={{\mathbb{R}}^{4d}}/H$ has the Lie algebra homomorphism $d\psi=d\eta=d\pi$, just like $\pi$. Here we use the fact that $H$ is discrete, thus $dq_{H}$ is the identity map.
Uniqueness. Since the domain of $\pi$ and $\psi$ is simply connected, $\psi=\pi$, i.e.,
(4.21) $q_{H}\circ\eta=q_{H}\circ\Phi.$
Kernels. The kernel of $\psi$ is $\ker q_{H}\circ\eta=\ker q_{H}=H$, whereas, by Lemma 4.4, $\ker q_{H}\circ\Phi=\{0\}$. Therefore, $H=\{0\}$ and $q_{H}$ is the identity map. Finally,
(4.22) $\Phi(\lambda)=\eta(\lambda)=\mathcal{A}\lambda,\quad\lambda\in{{\mathbb{R}}^{4d}}.$
This concludes the proof. ∎
It remains to prove the main result.
Proof of Theorem 1.1.
By Lemma 3.2, there exists a weak*-continuous operator ${\mathfrak{T}}:{\mathcal{S}({{\mathbb{R}}^{2d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ and an invertible matrix $\mathcal{A}\in{\mathrm{Sp}(4d,{\mathbb{R}})}$ such that for all $\lambda\in{{\mathbb{R}}^{4d}}$ and all $F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}$ holds
(4.23) ${\mathfrak{T}}\rho(\lambda)F=c\,\rho(\mathcal{A}\lambda){\mathfrak{T}}F.$
By Lemma 4.2, $\mathcal{A}$ satisfies
(4.24) $[\lambda,\mu]-[\mathcal{A}\lambda,\mathcal{A}\mu]\in{\mathbb{Z}},\qquad\lambda,\mu\in{{\mathbb{R}}^{4d}}.$
Since the expression is continuous in $\lambda$ and $\mu$, this implies that
(4.25) $[\lambda,\mu]-[\mathcal{A}\lambda,\mathcal{A}\mu]=0,\qquad\lambda,\mu\in{{\mathbb{R}}^{4d}}.$
i.e., $\mathcal{A}$ is symplectic.
Define ${\mathfrak{T}}_{0}:{\mathcal{S}({{\mathbb{R}}^{2d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ as
${\mathfrak{T}}_{0}\coloneqq\hat{\mathcal{A}}^{-1}{\mathfrak{T}}\,.$
Then by (2.4), ${\mathfrak{T}}_{0}$ satisfies the identity
(4.26) ${\mathfrak{T}}_{0}\rho(\lambda)F=\hat{\mathcal{A}}^{-1}{\mathfrak{T}}\rho(\lambda)F=\hat{\mathcal{A}}^{-1}\rho(\mathcal{A}\lambda){\mathfrak{T}}F=\rho(\lambda)\hat{\mathcal{A}}^{-1}{\mathfrak{T}}F=\rho(\lambda){\mathfrak{T}}_{0}F,$
that is, ${\mathfrak{T}}_{0}$ commutes with all time-frequency shifts. This fact implies that ${\mathfrak{T}}_{0}$ is a multiple of the identity operator on ${\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$.
If ${\mathfrak{T}}_{0}$ were bounded on $L^{2}({\mathbb{R}}^{d})$, then this conclusion would follow from Schur’s lemma. On the level of distributions, one may argue as follows. Since ${\mathfrak{T}}_{0}$ commutes in particular with all time shifts $T_{x}=\rho(x,0)$, it is a convolution operator [11, p. 169], i.e., there exists a $K_{0}\in{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ such that
(4.27) ${\mathfrak{T}}_{0}F=K_{0}*F,\quad F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}$
holds in the distributional sense. Since ${\mathfrak{T}}_{0}$ also commutes with modulations,
(4.28) $K_{0}*F=M_{-\omega}M_{\omega}{\mathfrak{T}}_{0}F=M_{-\omega}{\mathfrak{T}}_{0}M_{\omega}F=M_{-\omega}(K_{0}*M_{\omega}F).$
Since $\langle M_{\omega}K_{0},G\rangle=\langle K_{0},M_{\omega}G\rangle$ and $\langle K_{0}*F,G\rangle=\langle K_{0},\tilde{F}*G\rangle$, where $\tilde{F}(t)=F(-t)$ denotes the reflection of $F$, the last equation can be rewritten as
$\langle K_{0},\tilde{F}*G\rangle$ $=\langle M_{\omega}(K_{0}*M_{-\omega}F),G\rangle=\langle K_{0}*M_{-\omega}F,M_{\omega}G\rangle$
$=\langle K_{0},(M_{-\omega}F)\,\tilde{}*(M_{\omega}G)\rangle=\langle K_{0},(M_{\omega}\tilde{F})*(M_{\omega}G)\rangle$
$=\langle K_{0},M_{\omega}(\tilde{F}*G)\rangle=\langle M_{\omega}K_{0},\tilde{F}*G\rangle.$
Now, ${\mathcal{S}({{\mathbb{R}}^{2d}})}*{\mathcal{S}({{\mathbb{R}}^{2d}})}$ spans a dense subspace of ${\mathcal{S}({{\mathbb{R}}^{2d}})}$, because it contains all Gaussian functions. Therefore it follows that $K_{0}=M_{\omega}K_{0}$ for all $\omega\in{{\mathbb{R}}^{2d}}$. This implies that $K_{0}$ is supported on $\{0\}$. The only distribution supported on $\{0\}$ which commutes with all frequency shifts is a multiple of the identity, i.e., there exists an $a\in{\mathbb{C}}$ such that
(4.29) $\hat{\mathcal{A}}^{-1}{\mathfrak{T}}F={\mathfrak{T}}_{0}F=aF,\qquad F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}.$
Equivalently, ${\mathfrak{T}}=a\hat{\mathcal{A}}$ on ${\mathcal{S}({{\mathbb{R}}^{d}})}$. This concludes the proof of the main theorem. ∎
It seems that a homomorphism from a locally compact group $G$ to a quotient $G/H$ with respect to a closed normal subgroup $H$ can be one-to-one only if $H=\{e\}$, but currently we do not have an argument for this more general result.
5. Final Remarks
To appreciate the proof above, we present some alternative arguments. They do not require assume the measurability of $\Phi$, but require additional assumptions on the bilinear form ${\mathfrak{R}}$. We note that in our formulation of Theorem 1.1 these assumptions appear as corollaries.
Eigendistributions. The following argument is based on easy facts about the eigen-distributions of time-frequency shifts. It is easy to see that an eigen-distribution $\rho(\lambda)F=F$ must be in $F\in{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}\setminus\bigcup_{p\leq\infty}L^{p}({{\mathbb{R}}^{2d}})$. Just write $\lambda=\mathcal{A}(x,0)$ for $\mathcal{A}\in{\mathrm{Sp}(2d,{\mathbb{R}})}$ and $\rho(\lambda)F=\hat{\mathcal{A}}\rho(x,0)\hat{\mathcal{A}}^{-1}F=F$, then $\hat{\mathcal{A}}^{-1}F$ is periodic, and periodic distributions cannot be in $L^{p}({{\mathbb{R}}^{2d}})$ for $p<\infty$. Using this observation, we sketch a different version of Proposition 4.5 under additional hypothesis.
Proposition 5.1.
Let ${\mathfrak{T}}:{\mathcal{S}({{\mathbb{R}}^{2d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ be a non-zero linear operator satisfying the covariance property ${\mathfrak{T}}\rho(\lambda)F=c\rho(\Phi(\lambda)){\mathfrak{T}}F$ for all $\lambda\in{{\mathbb{R}}^{4d}}$ and all $F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}$.
Assume in addition that
(i) either there exist $F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}$ with a non-zero image ${\mathfrak{T}}F\in L^{p}({{\mathbb{R}}^{2d}})$ for some $p<\infty$, or
(ii) $\{{\mathfrak{T}}F:F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}\}$ is a weak-dense subspace of ${\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$.
Then $\Phi$ is an additive map.
Note that we do not assume that $\phi$ is measurable. Assumption (i) is quite reasonable, because a time-frequency representation of Schwartz functions should be “nice”. Assumption (ii) is very strong, but difficult to verify. It actually follows as a corollary from Theorem 1.1 provided that we assume the measurability of $\phi$.
Sketch of proof.
By Lemma 4.2
$\rho(\Phi(\lambda_{1}+\lambda_{2})-\Phi(\lambda_{1})-\Phi(\lambda_{2})){\mathfrak{T}}F=c{\mathfrak{T}}F$
with some phase factor $c$. Thus ${\mathfrak{T}}F\neq 0$ is an eigen-distribution of a time-frequency shift. For ${\mathfrak{T}}F\in L^{p}({{\mathbb{R}}^{2d}})$ this is only possible, if the time-frequency shift is a multiple of the identity, whence $\Phi(\lambda_{1}+\lambda_{2})-\Phi(\lambda_{1})-\Phi(\lambda_{2})=0$ and thus $\Phi$ is additive.
If ${\mathfrak{T}}({\mathcal{S}({{\mathbb{R}}^{d}})}\otimes{\mathcal{S}({{\mathbb{R}}^{d}})})$ spans a weak-dense subspace of ${\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$, then again the time-frequency shift on the left-hand side must be a multiple of the identity, and $\Phi$ is additive. ∎
Uniqueness. Some arguments would simplify, if we knew that the distortion function $\Phi$ were unique.
Proposition 5.2.
Let ${\mathfrak{R}}:{\mathcal{S}({{\mathbb{R}}^{d}})}\times{\mathcal{S}({{\mathbb{R}}^{d}})}\to{\mathcal{S}^{\prime}({{\mathbb{R}}^{2d}})}$ be a non-zero, bilinear, separately weak*-continuous mapping satisfying the general covariance property  (1.4). Further, assume that $\Phi_{\mathfrak{R}}$ is unique. Then there exist $a\in{\mathbb{C}}$, $a\neq 0$, and $\mathcal{A}\in{\mathrm{Sp}(4d,{\mathbb{R}})}$ such that
(5.1) ${\mathfrak{R}}(f,g)=a\,\hat{\mathcal{A}}(f\otimes g),\qquad f,g\in{\mathcal{S}({{\mathbb{R}}^{d}})}.$
Proof sketch.
The only part where measurability plays a role in the proof above is related to Proposition 4.5. We indicate how to use uniqueness instead. The rest of the proof is the same as the proof of Theorem 1.1. A simple argument, which we omit, shows that the constant in (1.4) can be taken to be $c=1$.
Now for $\lambda\in{{\mathbb{R}}^{4d}}$, $n\in{\mathbb{N}}$, and $F\in{\mathcal{S}({{\mathbb{R}}^{2d}})}$ we have
$\rho(\Phi(n\lambda)){\mathfrak{T}}F$ ${\mathfrak{T}}\rho(n\lambda)F=\rho(\Phi(\lambda)){\mathfrak{T}}\phi\big{(}(n-1)\lambda\big{)}F=\dots$
$=\rho(\Phi(\lambda))^{n}{\mathfrak{T}}F=\rho\big{(}n(\Phi(\lambda)\big{)}{\mathfrak{T}}F$
By uniqueness of $\Phi$ we obtain
$\Phi(n\lambda)=n\Phi(\lambda)\,,$
which implies even that $\Phi$ is ${\mathbb{Q}}$-homogeneous.
We now apply (4.3), where the measurability of $\Phi$ is not used, and obtain for all $m,n\in{\mathbb{Z}}$
(5.2) $2^{m+n}([\lambda,\mu]-[\Phi(\lambda),\Phi(\mu)])=[2^{m}\lambda,2^{n}\mu]-[\Phi(2^{n}\lambda),\Phi(2^{n}\mu)]\in{\mathbb{Z}}.$
This is equivalent to
(5.3) $[\lambda,\mu]-[\Phi(\lambda),\Phi(\mu)]=0,$
which already implies that $\Phi$ is a linear symplectic map.
The rest of the proof is identical to the Proof of Theorem 1.1. ∎